OpenStudy (anonymous):

Simplify: (5x+6)/(x^2-1) - (1)/(x^2-1)

OpenStudy (ghazi):

just take the base in common and you'll get the solution

OpenStudy (anonymous):

so take out (x^2-1) ?

OpenStudy (ghazi):

yes...

OpenStudy (anonymous):

so what would it be then?

OpenStudy (ghazi):

now you'll have (5x+6-1)/(x^2-1)

OpenStudy (ghazi):

simplify it ...now

OpenStudy (ghazi):

and take out 5 in common from numerator and factorize the denominator

OpenStudy (anonymous):

i dont know what to do at all...

mathslover (mathslover):

\[\large{\frac{5x+6}{x^2-1}-\frac{1}{x^2-1}}\] \[\large{\frac{5x+6-1}{x^2-1}}\] right?

OpenStudy (anonymous):

yeah, but wht next?

mathslover (mathslover):

what is 6 - 1 ?

OpenStudy (anonymous):

would it be 5/x-1 ?

OpenStudy (anonymous):

5

mathslover (mathslover):

very good now : \[\large{\frac{5x+6-1}{x^2-1}}\] \[\large{\frac{5x+5}{x^2-1}}\]

OpenStudy (ghazi):

no it would be 5(x+1)/(x+1)(x-1) now you'll get 5/(x-1)

mathslover (mathslover):

\[\large{\frac{5(x+1)}{(x+1)(x-1)}}\] \[\large{\frac{5\cancel{(x+1)}^1}{\cancel{(a+1)}^1(a-1)}}\] \[\large{\frac{5}{x-1}}\]

mathslover (mathslover):

oops sorry i meant x-1 there not a-1

mathslover (mathslover):

so got it @simplybeyoutiful ?

OpenStudy (anonymous):

yes, thats what i thought it was thank you

mathslover (mathslover):

ur welcome